### v4

@@ Abstract
Masala-CHAI is the first fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists. It addresses a long-standing challenge in automating netlist generation for analog circuits within circuit design automation. Automating this workflow could accelerate the creation of finetuned LLMs for analog circuit design and verification. We identify key challenges in this automation and evaluate the multi-modal capabilities of state-of-the-art LLMs, particularly GPT-4, to address these issues. We propose a three-step workflow to overcome current limitations: labeling analog circuits, prompt tuning, and netlist verification. This approach aims to create an end-to-end SPICE netlist generator from circuit schematic images, tackling the long-standing hurdle of accurate netlist generation. Our framework demonstrates significant performance improvements, tested on approximately 2,100 schematics of varying complexity. We open-source this solution for community-driven development.
@@ I Introduction
Large Language Models (LLMs) have received significant attention due to their wide-ranging applications, from text summarization to code generation, and have a growing impact across various fields. For hardware design, LLMs have primarily demonstrated potential in the digital domain. This includes tasks such as Verilog code generation [1, 2, 3, 4], assertion generation [5, 6], bug fixing [7, 8], and electronic design automation (EDA) tool scripting [9, 10]. The success of these domain-tailored LLMs relies on access to large and high-quality datasets. For instance, 75K Verilog files from GitHub were used to train the Verigen code generation model [4]. Data scraped from Verilog textbooks was also used to improve the performance of LLMs for Verilog [4]. Textbooks have also proven useful in other domains, for instance, in systems biology [11] and in understanding protein interactions [12].
@@
Building on the success of LLMs in the digital domain, it is natural to explore their application in the analog domain, specifically, in the automated generation of analog circuits from natural language specifications. Analog circuits are described in SPICE (and its many variants), the industry-standard textual representation for simulating analog circuits. SPICE is a low-level description that defines the interconnections between analog components like resistors, capacitors, inductors, and transistors. Open-source datasets for SPICE, unfortunately, are very limited compared to Verilog code, creating a gap in research in this area [13, 14, 15]. However, analog circuit textbooks and research papers are plentiful and contain a trove of analog circuit diagrams, but these are usually in image (or figure) format. These images must then be manually converted to SPICE netlists, which is painstaking and time-consuming.
To address this issue, a recently proposed method, AMSNet [13], introduced a semi-automated way to create a SPICE dataset from analog circuit figures using the capabilities of multi-modal LLMs like GPT-4o that take both image and text inputs and produce both image and text outputs. However, AMSNet still requires manual annotations of nets to obtain accurate SPICE netlists. Even with manual annotations, AMSNet still fails to extract accurate netlists of complex circuits (the reasons for this are described in Section III).
@@
Contributions: Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated method for SPICE netlist generation from analog circuit schematics. We incorporate various techniques, including (i) fine-tuned object detectors to extract passive and active components, (ii) deep Hough transform priors to identify nets, (iii) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (iv) post-extraction verification. We make four contributions:
@@
    Build Masala-CHAI, the first fully automated framework using LLMs along with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics present in textbooks and papers.
@@
    Use Masala-CHAI to collect a dataset of $\sim$2,100 SPICE netlists (along with metadata like figure captions) from a sample textbook. We open-source our flow to the community to further research in this area [16].
@@
    Fine-tune GPT models using the Masala-CHAI dataset to demonstrate the potential use-case of the dataset in automated SPICE netlist generation from English prompts.
@@ II Related Work
LLMs have gained significant attention in chip design [17]. Significant progress has been made in improving Verilog code generation [1, 2, 3, 4], with studies introducing new methodologies that enhance the quality of generated Verilog code. These advances demonstrate how LLMs can streamline and improve digital hardware design workflows. In addition to code generation, prompt engineering has proven effective in chip design [18, 19, 20]. By using LLMs, researchers have conceptualized and designed complex digital hardware architectures efficiently, facilitating faster and more accurate chip development. Beyond code generation, LLMs have found applications in assistant chatbots, script generation, and bug analysis [10]. Similarly, [9] explores LLMs in planning and executing tasks in the Electronic Design Automation (EDA) flow. The use of LLMs in generating assertions and testbenches for verifying the correctness of Integrated Circuit (IC) designs has also seen notable improvements [5, 6, 8, 7].
[21, 22] have explored schematic-to-netlist generation using ML-based approaches. However, these solutions are not reproducible due to the unavailability of any open-source implementation. Recent advancements have extended LLM applications into analog design. Closed-loop Python code generation for analog circuits using LLM agents has shown promising results [14]. LLMs have also demonstrated the ability to generate various circuit topologies from given specifications [23]. Furthermore, a dataset for exploring SPICE netlist generation was recently released in [13]. However, existing approaches require manual modifications of circuit schematics, which significantly limits their scalability for large-scale applications.
@@ III Key Challenges in SPICE Netlist Extraction
Despite advances in multi-modal LLM, as demonstrated by AMSNet [13], current solutions like GPT-4o do not automatically extract accurate SPICE netlists from circuit schematics. Our work begins with an in-depth analysis of the specific failure modes encountered by GPT-4o. Building on these insights, we propose solutions to overcome these limitations, advancing the field toward a fully automated SPICE netlist extraction process that significantly reduces dependence on manual annotations.
@@ III-A Can GPT-4o detect electrical components accurately?
When generating a SPICE netlist, accurately identifying all components is essential for designers but challenging for language models. To evaluate GPT-4o’s ability to recognize electrical components in circuit schematics, we curated a set of analog schematics featuring resistors, capacitors, inductors, MOSFETs, and various sources from a popular analog textbook [24]. As illustrated in Fig. 1, GPT-4o often misclassifies components, confusing NMOS and PMOS transistors (left and middle schematics) and even omitting key elements such as PMOS transistors and current sources in some cases. However, GPT-4o demonstrates the ability to identify all components accurately in some cases (final schematic), revealing promise despite its inconsistencies.
Solution 1: Electrical Component Detection using Object Detection Networks - Recognizing the limitations of GPT-4o in this task, we propose a dedicated solution using a CNN-based object detection model, using the strong performance of models such as YOLO [25]. By training a CNN model specifically designed for schematic components, we significantly improve the accuracy in detecting and bounding all circuit elements, a method detailed in more detail in Section IV.
Solution 2: Prompt Tuning - We developed a targeted prompt enhancement strategy to refine LLM component differentiation, especially between NMOS and PMOS transistors. By emphasizing structural differences in prompts (see Fig. 2), we achieved notable improvements in accuracy. Our empirical findings support this approach in Section IV.
III-B Can GPT-4o Properly Connect Circuit Components?
@@
Once all components are detected, the next step is ensuring that the components are connected correctly so the resulting netlist accurately reflects the schematic. While GPT-4o successfully maps 2-terminal devices, our study uncovered several critical failure modes that compromise the accuracy of netlist generation. These failure modes are illustrated in Fig. 3. Firstly, GPT-4o incorrectly assumes that intersecting nets are connected, even when no connection exists. This error arises partly from inconsistencies in how different schematic notations define connectivity. Secondly, GPT-4o frequently mixes up MOSFET terminals (drain, gate, source). Despite the drain and source’s electrical equivalence, this misidentification disrupts the final netlist’s correctness. Finally, differential input and output voltage pairs pose significant challenges for GPT-4o. For example, in the left schematic (Fig. 3), while components were correctly identified, the source connections for (M1 and M2) were incorrect, and the gate voltage for (M3 and M4) was wrongly assumed to be ‘Vb1’. Similar errors were observed in the right schematic, where differential output was mishandled, and the drain of M5 was incorrectly mapped.
Solution 1: Prompt Tuning - We enhanced the input prompts by explicitly specifying critical design features such as differential pairs and diode-connected topologies. This approach improved GPT-4o’s ability to map terminals to their corresponding nets for simpler designs. However, the solution’s scalability was limited, as it struggled with larger schematics containing more complex net connections.
Solution 2: Automatic Net Annotation - Leveraging our knowledge of component locations, we developed a framework to annotate nets with unique identifiers. By incorporating these annotations into the prompt (see Section IV), GPT-4o’s performance in translating schematics to SPICE netlists improved significantly. This systematic annotation reduced ambiguities and enhanced the LLM’s ability to correctly interpret and connect components, particularly in complex schematics.
@@ IV Methodology
This section details the methodology developed in this study, addressing the significant challenges outlined in Section III. Specifically, the accurate generation of SPICE netlists remains a critical bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections. To begin the dataset creation flow, as shown in Fig. 4 (➊), it takes a document in PDF format. After that, it is subdivided into 2 parts: (I) Extraction/Summarization of text present in the document comprises of the description of schematic and Key Performance Indicators (KPI) like gain, linearity, bandwidth, noise, gain/phase margin, etc - This will be used as an additional context for fine-tuning LLMs, Section V-D, shown in Fig. 4 (➏ - ➐); (II) Automatic SPICE Netlist Generation - Masala-CHAI: This step starts with extraction of the schematics/image from the document, then goes through the three-step process: 1) Labeling Analog Circuit, 2) Prompt Tuning, and 3) SPICE netlist verification, to finally generate a SPICE netlist, shown in Fig. 4 (➋ - ➎).
@@ IV-A Labeling Analog Circuit
Fig. 4 (➋) step eases the job of the LLM in writing the SPICE netlist from schematics by detecting components and enabling net annotation. As discussed in Section III, SPICE netlist generation is not straightforward, requiring the designer to feed information alongside the schematics. However, this manual process hinders generating large-scale datasets. Thus, we automate labeling all parts of the schematics.
@@ IV-A1 Detection of Circuit Components
We train YoloV8 [25], a state-of-the-art object localization and classification model, specifically for detecting and classifying circuit components with high precision, similar to prior work [21, 22]. Given an input circuit image, YoloV8 localizes each component by regressing its center coordinates, bounding box parameters, and classification label. The model’s backbone, CSPDarknet53 [26], enhanced with a Cross Stage Partial bottleneck [27], effectively integrates high-level semantic information with low-level spatial details, enabling accurate detection even for small-scale components — a critical requirement for high-performance circuit analysis. This process results in schematic diagrams annotated with bounding boxes for all detected components, as illustrated in Fig. 5 (➋). We utilize YoloV8 on the open-sourced dataset [28]. This dataset comprises approximately 4,300 circuit diagram snapshots annotated with bounding boxes across 12 component classes: AC Source, BJT, Battery, Capacitor, DC Source, Diode, Ground, Inductor, MOSFET, Resistor, Current Source, Voltage Source. We resized images to 640x640 and trained for 1000 epochs with a learning rate of 0.01.
@@ IV-A2 Net Detection
We performed net detection in circuit schematics by leveraging pre-trained Deep Hough Transform Line priors [29], a state-of-the-art approach for line detection in images. Unlike conventional methods, this model operates in the Hough domain, parametrizing line segments in polar coordinates to achieve precise line segmentation. We remove components from the schematic to address challenges posed by line segments within components such as capacitors and MOSFETs, leaving only the lines representing nets, as illustrated in Fig. 5 (➌). This ensures a clean input for subsequent net analysis and eliminates potential sources of ambiguity. Following net detection, we propose a simple yet effective heuristic for clustering line segments into nets. Specifically, we group all line segments into a single cluster if their endpoints fall within a radius of 40 pixels, as demonstrated in Fig. 5 (➍). To ensure reliability, we fine-tune the radius parameter and manually verify the consistency of the clustered nets.
@@ IV-B Prompt Tuning
Prompts are crucial in guiding LLMs to generate precise, high-quality outputs. Crafting effective prompts involves meticulous design of their language, structure, and contextual framing, enabling models to respond accurately and meaningfully to user inputs. Through prompt refinement, designers can precisely control model behavior, ensuring the generation of coherent and domain-specific responses. In this study, we systematically explore diverse prompt designs tailored to enhance understanding of schematics. Our approach demonstrates significant effectiveness, as illustrated in Fig. 2 and Fig. 6. Specifically, prompt in Fig. 6 is used at the stage Fig. 4 (➌), facilitating insightful queries to the GPT model.
You are an expert analog designer. You will be provided with a schematic, your task is to follow the below instructions carefully: 1. List all the components which you can observe from the figure. 2. MOSFET are 3 terminal devices with (drain, gate, source). 3. For each component, look at the net number highlighted in red. 4. To identify the source terminal of a MOSFET, choose the net highlighted in red which is nearest to the arrow of the MOSFET. 5. Write a SPICE netlist.
Figure 6: Prompt Tuning guides the GPT to generate SPICE netlist from Schematics.

### v5

@@ Abstract
Masala-CHAI is a fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists. It addresses a long-standing challenge in circuit design automation: automating netlist generation for analog circuits. Automating this workflow could accelerate the creation of fine-tuned LLMs for analog circuit design and verification. In this work, we identify key challenges in automated netlist generation and evaluate multimodal capabilities of state-of-the-art LLMs, particularly GPT-4, in addressing them. We propose a three-step workflow to overcome existing limitations: labeling analog circuits, prompt tuning, and netlist verification. This approach enables end-to-end SPICE netlist generation from circuit schematic images, tackling the persistent challenge of accurate netlist generation. We utilize Masala-CHAI to collect a corpus of 7,500 schematics that span varying complexities in 10 textbooks and benchmark various open source and proprietary LLMs. Models fine-tuned on Masala-CHAI when used in LLM-agentic frameworks such as AnalogCoder achieve a notable 46% improvement in Pass@1 scores. We open-source our dataset and code for community-driven development.
@@ I Introduction
Large Language Models (LLMs) have received significant attention due to their wide-ranging applications, from text summarization to code generation, and have a growing impact across various fields. For hardware design, LLMs have primarily demonstrated potential in the digital domain. This includes tasks such as Verilog code generation [1, 2, 3, 4], assertion generation [5, 6], bug fixing [7, 8], and electronic design automation (EDA) tool scripting [9, 10]. The success of these domain-tailored LLMs relies on access to large and high-quality datasets. For example, the Verigen code generation model [4] was trained on a dataset of 75,000 Verilog files sourced from GitHub, with additional data extracted from Verilog textbooks to further enhance its performance in Verilog-related tasks. Beyond hardware design, textbooks have also been valuable in other domains, such as systems biology [11] and protein interaction studies [12], demonstrating their broader utility in training domain-specific models.
Figure 1: Pass@1 performance of GPT-4o fine-tuned with Masala CHAI datasets extracted from between 1-10 textbooks, compared to the no fine-tuning baseline (task IDs from Table II). Our largest dataset of  7500 captioned SPICE netlists extracted from 10 textbooks provides more than 40% Pass@1 accuracy even on challenging tasks like telescoping cascode amplifier and bandgap reference circuit generation.
@@
Building on the success of LLMs in the digital domain, it is natural to extend their application in the analog domain, particularly in the automated generation of analog circuits from natural language specifications. Analog circuits are described in SPICE and its various variants, which serve as the industry-standard textual representation for analog circuit simulation. SPICE provides a low-level description that defines the interconnections between analog components like resistors, capacitors, inductors, and transistors. The availability of open-source datasets for SPICE remains significantly limited compared to Verilog, which poses a major challenge for automating the learning of frameworks for the generation of net lists [13, 14, 15, 16, 17]. However, analog circuit textbooks and research papers are plentiful and contain a trove of analog circuit diagrams, but these are usually in image (or figure) format. These images must then be manually converted to SPICE netlists, which is painstaking and time-consuming.
@@
Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated framework for SPICE netlist generation from analog circuit schematics. We incorporate various techniques, including (i) Schematic extractor from documents, (ii) fine-tuned circuit component detectors to extract passive and active components, (iii) deep hough transform priors to identify nets, (iv) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (v) post-extraction verification. We curated the largest open-source analog circuit dataset, comprising 7,500 examples using Masala-CHAI, and fine-tuned various LLMs, including CodeLlama, DeepSeek-34B, and GPT-(3.5,4). As shown in  Figure 1, larger datasets significantly enhance performance, improving pass@k results across different numbers of textbooks used for fine-tuning.
Our contributions are as follows:
@@
    Build Masala-CHAI, the first fully automated framework that combines LLMs with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics found in textbooks and research papers.
@@
    Leverage Masala-CHAI to compile an open-sourced dataset of $\sim$7,500 SPICE netlists with metadata (figure captions, components, and nodes) from 10 textbooks. To the best of our knowledge, this is the largest curated and labeled SPICE dataset available currently.
@@
    Finetune and benchmark four state-of-art LLMs (CodeLLama-70B, DeepSeek-34B, GPT-3.5, GPT-4o) on the Masala-CHAI dataset, showcasing its potential for automated SPICE netlist generation from English prompts.
  5. 5.
[…]
@@ II Related Work
LLMs have gained significant attention in chip design [18]. Tremendous progress has been made in improving Verilog code generation [1, 2, 3, 4]. These advances demonstrate how LLMs can streamline and improve digital hardware design workflows. In addition to code generation, prompt engineering has proven effective in chip design [19, 20, 21]. Using LLMs, researchers have efficiently conceptualized and designed complex digital hardware architectures, facilitating faster and more accurate chip development. Beyond code generation, LLMs have found applications in assistive chatbots, script generation, and bug analysis [10].  [9] explores LLMs in planning and executing tasks in the Electronic Design Automation (EDA) flow. The use of LLMs in generating assertions and testbenches for verifying the correctness of Integrated Circuit (IC) designs has also seen notable improvements [5, 6, 8, 7].  [16, 17] have explored schematic-to-netlist generation using ML-based approaches. However, these solutions are not reproducible due to the unavailability of any open-source implementation. Recent advances extended LLM applications into analog design. Closed-loop Python code generation for analog circuits by LLM agents has shown promising results [14]. LLMs are able to generate circuit topologies from specifications [22]. A dataset for exploring SPICE netlist generation was recently released in [13], summarized in Table I.
TABLE I: Comparison with AMSNet [13]. Here, ‘M’ denotes Manual, ‘P’ denotes Partial Automation (with some manual intervention), and ‘A’ denotes Fully Automated.
Schematic Component Nets Textbooks Dataset
AMSNet [13] M A P 1 734
Masala-CHAI A A A 10 7500
@@ III Key Challenges in SPICE Netlist Extraction
Despite advances in multi-modal LLMs, models like GPT-4o do not automatically extract accurate SPICE netlists from circuit schematics [13]. Our work begins with an in-depth analysis of GPT-4o’s failure modes. Building on these insights, we propose our solutions and introduce a fully automated SPICE netlist extraction framework without manual annotations.
@@ III-A Can GPT-4o detect electrical components accurately?
When generating a SPICE netlist, accurately identifying all circuit components is essential for designers and challenging for language models. We curated a set of analog schematics featuring resistors, capacitors, inductors, MOSFETs, and various sources from a popular analog textbook [23] and evaluated GPT-4o’s ability to recognize electrical components in circuit schematics. As shown in Figure 2, GPT-4o frequently misidentifies components, often confusing NMOS and PMOS transistors (left and middle schematics) or omitting key elements like PMOS transistors and current sources. However, in some cases (final schematic), it accurately identifies all components, demonstrating potential despite its inconsistencies.
Solution 1: Customized Object Detection Network for Analog Circuit Components- Recognizing the limitations of GPT-4o in this task, we propose a dedicated solution using a deep CNN-based object detection model, using the strong performance of models such as YOLO [24]. By training a CNN model designed explicitly for schematic components, we significantly improve the accuracy in detecting and bounding all circuit elements, a method detailed in more detail in Section IV.
Solution 2: Prompt Tuning - We developed a targeted prompt enhancement strategy to refine LLM component differentiation, especially between NMOS and PMOS transistors. By emphasizing structural differences in prompts (see Figure 3), we achieved notable improvements in accuracy. Our empirical findings support this approach in Section IV.
III-B Can GPT-4o accurately connect circuit components?
@@
After detecting all circuit components, the next step is to ensure their correct connections, so the generated netlist accurately represents the schematic. While GPT-4o successfully maps 2-terminal devices, our study uncovered several critical failure modes that compromise the accuracy of netlist generation. These failure modes are illustrated in Figure 4. Firstly, GPT-4o incorrectly assumes that intersecting nets are connected, even when no connection exists. This error arises partly from inconsistencies in the definition of connectivity across different schematic notations. Secondly, GPT-4o often misidentifies MOSFET terminals, confusing the drain, gate, and source. Despite the drain and source’s electrical equivalence, this misidentification affects the accuracy of the final netlist. Finally, differential input and output voltage pairs pose significant challenges for GPT-4o. For example, in the left schematic of Figure 4, while components were correctly identified, the source connections for (M1 and M2) were incorrect, and the gate voltage for (M3 and M4) was wrongly assumed to be ‘Vb1’. Similar errors appeared in the right schematic, where differential output was mishandled, and M5’s drain was incorrectly mapped.
Solution: Automatic Net Annotation using Deep Hough Transform - Leveraging our knowledge of component locations, we developed a method to identify and annotating nets using a Deep Hough transform line priors. Incorporating these annotations into the prompt (see Section IV), led to improvement in GPT-4o’s performance in translating schematics to SPICE netlists. This systematic annotation reduced ambiguities and enhanced the LLM’s ability to interpret and connect components, particularly in complex schematics correctly.
@@ IV Methodology
This section details the methodology used, addressing the challenges outlined in Section III. Accurate generation of SPICE netlists remains a bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections. Our dataset creation framework, shown in Figure 5, starts by paginating a textbook into individual PDFs (➊), which are then passed through a YOLOv8 [24], a state-of-the-art object localization and classification model, to extract schematics from the page. Along with extracted schematics, we also extract and summarize text describing the schematic, for example, from the schematic caption or text on the page referring to the schematic. These are used as additional context for fine-tuning LLMs, Section V-B, shown in Figure 5 (➐ - ➒). The extracted schematic is still in image format and needs to be reliably converted to a SPICE netlist. To this end, Masala-CHAI follows a three-step process: (i) labeling analog circuits, (ii) prompt tuning, and (iii) SPICE netlist verification, starting from schematic extraction ( Figure 5, ➌ - ➏). These steps are described next.
@@ IV-A Labeling Analog Circuit
Figure 5 (➌) simplifies the LLM’s task of generating SPICE netlists by detecting components and facilitating net annotation. As discussed in Section III, SPICE netlist generation is not straightforward, requiring the designer to feed information alongside the schematics. However, this manual process hinders generating large-scale datasets. Thus, we automate the labeling of all parts of the schematics.
@@ IV-A1 Detection of Circuit Components
We train a custom YoloV8 model [24] specifically for detecting and classifying circuit components with high precision, similar to prior work [16, 17]. YoloV8’s Darknet53 backbone [25], enhanced with a “Cross Stage Partial” bottleneck [26], integrates high-level semantic information with low-level spatial details, enabling accurate detection even for small components—a requirement for component detection in complex circuit schematics. We fine-tune YOLOv8 on an open-source dataset [27] consisting of  4,300 circuit diagrams with annotated bounding boxes across 12 component classes: AC Source, BJT, Battery, Capacitor, DC Source, Diode, Ground, Inductor, MOSFET, Resistor, Current Source, and Voltage Source. Images were resized to 640x640 and trained for 1000 epochs with a learning rate of 0.01. The resulting fine-tuned YoloV8 model localizes each component in the schematic by outputting its center coordinates, a bounding box, and classification labels, as illustrated in Figure 6 (➋).
@@ IV-A2 Net Detection
We performed net detection in circuit schematics by leveraging pre-trained deep hough transform line priors [28], a state-of-the-art approach for line detection in images. Unlike conventional methods, this model operates in the Hough domain, parametrizing line segments in polar coordinates to achieve precise line segmentation. We remove components from the schematic to address challenges posed by line segments within components such as capacitors and MOSFETs, leaving only the lines representing nets, as illustrated in Figure 6 (➌). This ensures a clean input for subsequent net analysis and eliminates potential sources of ambiguity.
Following net detection, we propose a simple yet effective heuristic for clustering line segments into nets. Specifically, we group all line segments into a single cluster if their endpoints fall within a radius of 40 pixels, as demonstrated in Figure 6 (➍). To ensure reliability, we tune the radius parameter and manually verify the consistency of the clustered nets.
@@ IV-B Prompt Tuning
Prompts are crucial in guiding LLMs to generate precise, high-quality outputs. Crafting effective prompts involves meticulous design of their language, structure, and contextual framing, enabling models to respond accurately and meaningfully to user inputs. Through prompt refinement, designers can precisely control model behavior, ensuring the generation of coherent and domain-specific responses. In this study, we systematically explore diverse prompt designs tailored to enhance understanding of schematics. Our approach demonstrates significant effectiveness, as illustrated in Figure 3 and Figure 7. Specifically, prompt in Figure 7 is used at the stage Figure 5 (➍), facilitating insightful queries to the GPT model.
You are an expert analog designer. You will be provided with a schematic, your task is to follow the below instructions carefully: 1. List all components which you can observe from the figure. 2. MOSFETs are 3 terminal devices with (drain, gate, source). 3. For each component, look at the net number in red. 4. To identify source terminal of a MOSFET, choose net highlighted in red which is nearest to the arrow of the MOSFET. 5. Write a SPICE netlist.
Figure 7: Prompt GPT to generate SPICE netlist from Schematic.